OpenStudy (anonymous):

1. In this problem, we will analyze the profit found for sales of decorative tiles. A demand equation (sometimes called a demand curve) shows how much money people would pay for a product depending on how much of that product is available on the open market. Often, the demand equation is found empirically (through experiment, or market research). a. Suppose that a market research company finds that at a price of p = $40, they would sell x = 40 tiles each month. If they lower the price to p = $30, then more people would purchase the tile, and they can expect to sell x = 50 tiles in a month’s

OpenStudy (anonymous):

Sorry for some reason the end of the question cut off. Here it is. sell x = 50 tiles in a month’s time. Find the equation of the line for the demand equation. Write your answer in the form p = mx + b. (Hint: Write an equation using two points in the form (x,p)).

OpenStudy (anonymous):

I am completely lost here. Somebody please help.

OpenStudy (anonymous):

p-p1=m(x-x1) p-40=(40-30/40-50)(x-40) p-40=-1(x-40) p=-x+80

OpenStudy (anonymous):

Ok. Why does p=p-p1 and why does m=-1 and why does b=-x1

OpenStudy (anonymous):

i used the point slope formula p does not equal p-p1

OpenStudy (anonymous):

but the question says to write the equation in the form p=mx+b and you have p-p1=m(x-x1)

OpenStudy (anonymous):

right, i then used algebra to move around the information so that it fits p=mx+b m=-1 and b=80

OpenStudy (anonymous):

ok